OLIVER69:

1. If you roll a die twice, what is the probability at least one of the die shows a 5? 2. If you roll a die twice, what is the probability the you will roll a sum of 6? 3. A product code consists of three letters chosen from the set {A, D, T, G}. The letters can be repeated. What is the probability a code contains exactly one D? 4. product code consists of three letters chosen from the set {L, M, N, O}. The letters cannot be repeated. What is the probability a code contains an L?

seashellpower:

AFRICA

seashellpower:

i think

seashellpower:

@luhivqqcherry

luhivqqcherry:

africa is insane .

seashellpower:

am i right??????

luhivqqcherry:

no .

luhivqqcherry:

girl u is slow .

seashellpower:

ohh..

luhivqqcherry:

LOL .

seashellpower:

ur a bully

luhivqqcherry:

no .

seashellpower:

im telling james

luhivqqcherry:

smh .

luhivqqcherry:

now let me answer olivers question .

seashellpower:

ur gonna get it wrong

seashellpower:

dont google yonna

seashellpower:

๐Ÿ˜ˆ

luhivqqcherry:

wow .

luhivqqcherry:

1. 11/36 . 2. 5/36 . 3. 3/4 . 4. 3/4 .

Markthegoat1117:

1. To calculate the probability of at least one die showing a 5 when rolling a die twice, we can use the concept of complementary probability. The probability of not rolling a 5 on a single die roll is 5/6, as there are 6 possible outcomes and only 1 of them is a 5. Since the two die rolls are independent events, the probability of not rolling a 5 on either roll is (5/6) * (5/6) = 25/36. Therefore, the probability of at least one die showing a 5 is 1 - 25/36 = 11/36. 2. To calculate the probability of rolling a sum of 6 when rolling a die twice, we need to determine the number of favorable outcomes and the total number of possible outcomes. The favorable outcomes are when the two dice show numbers that sum up to 6. These outcomes are: (1, 5), (2, 4), (3, 3), (4, 2), and (5, 1). So, there are 5 favorable outcomes. The total number of possible outcomes when rolling two dice is 6 * 6 = 36, as each die has 6 possible outcomes. Therefore, the probability of rolling a sum of 6 is 5/36. 3. To calculate the probability of a product code containing exactly one D, we need to consider the total number of possible codes and the number of codes that meet the given condition. Since the code consists of three letters chosen from the set {A, D, T, G}, there are 4 choices for each letter. Therefore, the total number of possible codes is 4 * 4 * 4 = 64. To have exactly one D in the code, we can place the D in any of the three positions (first, second, or third). The remaining two positions can be filled with any of the other three letters (A, T, G). So, there are 3 * 3 * 3 = 27 codes that meet the given condition. Therefore, the probability of a code containing exactly one D is 27/64. 4. To calculate the probability of a product code containing an L when the letters cannot be repeated, we need to consider the total number of possible codes and the number of codes that contain an L. Since the code consists of three letters chosen from the set {L, M, N, O}, there are 4 choices for the first letter, 3 choices for the second letter (excluding the one already chosen), and 2 choices for the third letter (excluding the two already chosen). Therefore, the total number of possible codes is 4 * 3 * 2 = 24. To have an L in the code, we can place the L in any of the three positions. The remaining two positions can be filled with any of the other three letters (M, N, O). So, there are 3 * 3 * 3 = 27 codes that contain an L. Therefore, the probability of a code containing an L is 27/24, which simplifies to 9/8.

OLIVER69:

Thanks @markthegoat1117 for the wonderful explanation to help me understand more. I think I got it now.

Markthegoat1117:

@oliver69 wrote:
Thanks @markthegoat1117 for the wonderful explanation to help me understand more. I think I got it now.
Glad to help๐Ÿ˜